Answer please
Arrange the following liquids in order of increasi
Answer #1

Whenever we add solid solute to a liquid solvent it's freezing point decrease which is called Freezing point depression.

Freezing point depression is a colligative property which depend on moles of solute particles.

∆Tf=iKf.m where Kf is constant. So the value of Freezing point depression depend on product of i(vant hoff factor) and m(molality).

So order of Freezing point would be:-(lowest to highest)aq MgI2(i×m=0.66) aq NaF(i×m=0.62), aq glucose (i×m=0.6), aq sucrose (i×m=0.5), H2O